OpenStudy (anonymous):

given sin A=3/5 and tan=-3/4. A and B are in QII. Find values of tan (A+B)

OpenStudy (anonymous):

@ajoek tan(??) = -3/4

OpenStudy (anonymous):

Ohh sorry...i mean...sin B=-3/4

OpenStudy (anonymous):

ok by using the formula tan(A+B) = tanA+tanB divided 1 - tanA*tanB

OpenStudy (anonymous):

now put tanA=sinA/cosA and tanB= sinB/cosB

OpenStudy (anonymous):

u can find the values of cosines by using the formula Sin^2A+Cos^2A=1

OpenStudy (anonymous):

@ajoek right??

OpenStudy (anonymous):

yes, agree